OpenStudy (anonymous):

Help! Given f (x) = 3x^3 - 4x^2 + x-1 what is f(2) Factor 36n^2 - 81 Simplify : mm(m^5)^0 m^-3 divided by m^6(y^3)^-3 solve : y^2 = 18

OpenStudy (lasttccasey):

So there are four parts to solve here? Given f (x) = 3x^3 - 4x^2 + x-1 what is f(2) Factor 36n^2 - 81 Simplify : mm(m^5)^0 m^-3 divided by m^6(y^3)^-3 Solve : y^2 = 18

OpenStudy (anonymous):

no they four differen questions

OpenStudy (anonymous):

no they are four different questions

OpenStudy (anonymous):

For the first one you plug in the 2 for x and solve that.

OpenStudy (lasttccasey):

Yes that's what I mean. Well the first one is fairly simple. Given f (x) = 3x^3 - 4x^2 + x-1 what is f(2) f(2) = 3(2)^3 - 4(2)^2 + (2) - 1 Basically just plugging in 2 for all 'x' variables f(2) = ... continue by doing the math and you should get your answer

OpenStudy (anonymous):

Go ahead lasttccasey. No point in two helping one person haha. I'll stick around though.

OpenStudy (lasttccasey):

Factor 36n^2 - 81 Since both have perfect squares you can unfoil them. Somewhat hard to explain but the format will be (x + (#))(x + (#)) where the # can be + or - Are you familiar with unfoiling?

OpenStudy (lasttccasey):

Sorry damain94 I didn't realize anyone answered before I hit post.

OpenStudy (anonymous):

It's fine your explanations and format are easier to understand.

OpenStudy (anonymous):

yes I am familiar with it

OpenStudy (lasttccasey):

Okay and for the third one, Simplify : mm(m^5)^0 m^-3 divided by m^6(y^3)^-3 What is mm? is that supposed to be m*m = m^2 or is mm another variable by itself?

OpenStudy (anonymous):

m*m

OpenStudy (lasttccasey):

Simplify : (m*m(m^5)^0 m^-3) / (m^6(y^3)^-3) Well first I'd multiply everything through. (m*m(m^5)^0 m^-3) / (m^6(y^3)^-3) becomes (m^-1) / (m^6*y^-9) any negative exponents are basically 1/ that term. ex. x^-2 = 1/x^2 (1/m)/(m^6/y^9) used the above rule y^9/m^7 multiplied through to get final answer.

OpenStudy (lasttccasey):

solve : y^2 = 18 Basically just square root both sides. y = + or - sqrt(18)